

********************************************************************
UARS WEEKLY STATUS REPORT - 02 MAR 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
SUSIM performed its daily solar scans of the sun this past week during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
PEM/AXIS-1 unit was activated (in half power mode) between February 12 and
February 16. At the present time, energy deposition from PEM/AXIS is not being
computed due to a software problem which PEM is to repair.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine. GSFC is
helping PEM with some DEC Alpha particulars which have caused PEM level 3
code to fail on DEC Alpha machines (software below level 3 executes on the
DEC Alpha).
The PEM team is working on replacing the PEM:: machine with a DEC Alpha
machine to be called PEM2K::. Initial tests of telemetry transfers through
PEM2K:: have been successful; however, the source of these failures has been
traced to quota limits and these have been corrected. PEM Telemetry is runing
in duplicate streams through PEM:: and PEM2K::. We find that files which
transfer from the CDHF to PEM:: and PEM2K:: are now identical. PEM has
requested DAP command capability with the CMS from PEM2K::. Shortly, PEM
will test the NRT capabilities of the rehost system, PEM::, and PEM2K::.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
HRDI continues to function normally, making wind measurements during the sunlit
parts of orbits nearly every day, and measuring nighttime winds when power is
available.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Daily processing of instrument Daily Activity Plans continues
with no problems. Instrument operations schedule is now complete
to 19 April 2001.
***********
CDHF
***********
DAAC DATA:
Data created in the month of December, 2000 for HALOE, NMC and UKMO has
been loaded into the DAAC Transfer Tables. DAAC has been notified.
UARS CDHF REHOST:
- Confirming all NRT users and that these users have access to NRT on
the Rehost. After Confirmation we will move NRT from current system to
Rehost. Also since we are ingesting l0 on both the rehost and current
system, those users who automatically receive L0 should be receiving 2
copies.
- The following software has been recompiled and relinked and PC is
testing new executables. Status follows:
- Attitude Extraction V7.3 successful and promoted to production mode.
Output test files were promoted from _TEST to _PROD. Being processed
on Current and Rehost Systems during (2) week parallel processing.
- Sensor Extraction V5.0 successful and promoted to production mode.
Being processed on Rehost System only per FDF.
- Observatory Log V3.4 successful and promoted to production mode.
Output test files were promoted from _TEST to _PROD. Being processed
on Rehost System only.
- HALOE V4.82 successful. Currently waiting validation results from PI.
- HRDI V4.4 successful. Currently waiting validation results from PI.
- SUSIM V7.2 testing in progress.
***********
FOT
***********
REPORTING PERIOD
This report covers 17 February 2001 (Orbit 51586, GMT Day 048) through
23 February 2001 (Orbit 51690, GMT Day 054).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing)
and is performing nominally under two-battery operations. The
instrument operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
-None this report period
The beta angle ranged between the angles of 28.7 degrees to 45.3 degrees
for this report period. The beta angle is now increasing towards a
maximum angle of 48 degrees on 26 February 2001 (DOY 057).
Spacecraft battery 1, 2, and 3 performance monitoring continues.
Battery 1 remains OFF the charge relay. Battery 2 and Battery 3 are
currently maintaining greater than 23.8 V end-of-night (EON) load bus
voltage. Battery temperatures are stable with a temperature delta
between Battery 2 and Battery 3 of 0.5 and 0.8 degrees C. Additional
battery performance data for each battery on 16 February 2001 and 23
February 2001 are provided in Appendix D.
The clock error ranged from -9.8 to 10.2 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and
Out-Of-Limits occurrences are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed
operational
element(s) during this report period (see Appendix A).
¥ FOT = 19
¥ SSPP = 1
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
¥ FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period
(see Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/027 - 2287.499274
GROUND SYSTEM OPERATION
One (1) operational support problem occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 01/053 show the observatory data loss to
be 300 hours, 19 minutes, 59.848 seconds. This shows no change since the
last report period This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the
mission.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
UARS is using both the Fine Sun Sensor (FSS) and the Three Axis
Magnetometer (TAM) as inputs to UFLTR this week. The FOT is pleased
with their performance, but continues to make adjustments to the TAM
alignment to improve its performance and UARS' attitude.
MEETINGS
The FOT attended a UARS SERB Meeting on Wednesday, February 21, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 15 March 2001 (DOY 074). This
maneuver will be a Forward to Reverse Flight maneuver (Northern
Hemisphere viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 16 February 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
51590 048/0655 01-105 NBTR NOOPS D-048
51597 048/1818 01-108 NBTR NOOPS D-049
51604 049/0618 01-108 NBTR NOOPS D-049
51612 049/1819 01-109 NBTR NOOPS D-050
51619 050/0619 01-109 NBTR NOOPS D-050
51627 050/1820 01-110 NBTR NOOPS D-051
51628 050/2042 01-111 OBC/TAM MAGMOD UPDATE
51634 051/0620 01-110 NBTR NOOPS D-051
51636 051/0842 01-112 OBC Clk Adj from 0.0 to 20.6 msec/day
51641 051/1640 01-114 OBC/TAM Minor TAM Alignment
51642 051/1821 01-113 NBTR NOOPS D-052
51649 052/0621 01-113 NBTR NOOPS D-052
51657 052/1822 01-115 NBTR NOOPS D-053
51659 052/2224 01-116 OBC Clk Adj from 20.6 to 0.0 msec/day
51664 053/0623 01-115 NBTR NOOPS D-053
51667 053/1127 01-115 NBTR NOOPS D-053
51672 053/1822 01-117 NBTR NOOPS D-054
51672 053/1822 01-118 SSPP Star Catalog
51680 054/0705 01-117 NBTR NOOPS D-054
51687 054/1823 01-119 NBTR NOOPS D-055
________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 23 February 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 23 February 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
51636 051/0844 Clk Adj from 0.0 to 20.6 msec/day
51659 052/2228 Clk Adj from 20.6 to 0.0 msec/day
________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 23 February 2001
16 Feb 2001 - GMT Day 047 Beta = 25.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ----------- ------- ---------- ----- ------- -----
1 +000.0/-022.4 -3.05/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +22.4/-67.2 +5.47/+4.19 +29.6/-14.4 25.4 1.040 20.6
3 +28.0/-44.8 +4.51/+2.92 +29.2/-14.4 25.4 1.030 20.5
23 Feb 2001 - GMT Day 054 Beta = 45.3 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ---------- -------- --------- ----- ------- -----
1 +000.0/-22.4 -3.05/-4.62 +0.0/+0.0 13.0 0.000 0.0
2 +11.2/-22.4 +4.51/+3.23 +27.2/-13.6 26.4 1.030 19.5
3 +22.4/-22.4 +3.87/+2.60 +26.4/-13.6 26.2 1.030 19.5
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
---- ---- ---------
-no changes V/T 5
________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 23 February 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-012 FOT 51591 048/0937 23619 LATE ACQUISITION
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-012 FOT 51591 048/0937 23619 LATE ACQUISITION
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-009 FOT see list Feb 01 14170 Generic Late
Acquisition
SEVNTFW01/08